# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 4.4 Standard Gaussian random design matrix
Define the event $\mathcal{E}(\mathbf{X}_{\mathcal{A}})\triangleq\{\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}\geq 9/(n\gamma_{\min}(\mathbf{\Sigma}))\}$. Since $|\mathcal{A}|\leq n$, the event $\mathcal{E}(\mathbf{X}_{\mathcal{A}})$ happens with probability at most $2e^{-n/2}$. Therefore, with probability at least $1-2e^{-n/2}$, the quantity $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter at most $9|\mathcal{A}|/(n\gamma_{\min}(\mathbf{\Sigma}))$. A Chernoff bound yields
$\mathbb{P}\left[\max_{j\in\mathcal{A}^{c}}|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\geq\frac{\bar{\alpha}}{2}\right]$ $\leq\mathbb{P}\left[\max_{j\in\mathcal{A}^{c}}|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\geq\frac{\bar{\alpha}}{2}~{}|~{}\mathcal{E}^{c}(\mathbf{X}_{\mathcal{A}})\right]+\mathbb{P}[\mathcal{E}(\mathbf{X}_{\mathcal{A}})]$
$\leq 2(d-|\mathcal{A}|)\exp\left(-\frac{\bar{\alpha}^{2}n\gamma_{\min}(\mathbf{\Sigma})}{72|\mathcal{A}|}\right)+2\exp(-n/2)$
$\leq 4(d-|\mathcal{A}|)\exp\left(-\frac{\bar{\alpha}^{2}n\gamma_{\min}(\mathbf{\Sigma})}{72|\mathcal{A}|}\right).$
@@
Therefore, with high probability, $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\leq 1-\bar{\alpha}/2$ for any $\mathbf{u}\in\mathbb{R}^{|\mathcal{A}|}$ such that $\|\mathbf{u}\|_{\infty}=1$. Summing over all $\mathcal{A}\subseteq[d]$ such that $|\mathcal{A}|\leq T$, we conclude that
$\mathbb{P}\left[\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}^{c}}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{\infty}\geq 1-\frac{\bar{\alpha}}{2}\right]$ $\leq 4d\exp\left(-\frac{\bar{\alpha}^{2}n\gamma_{\min}(\mathbf{\Sigma})}{72T}\right)\sum_{i=0}^{T}\binom{d}{i}$
$\leq 4d\exp\left(-\frac{\bar{\alpha}^{2}n\gamma_{\min}(\mathbf{\Sigma})}{72T}+d\mathsf{H}(T/d)\right),$
where $\mathsf{H}(p)\triangleq-p\ln{p}-(1-p)\ln(1-p)$ is the entropy function. Using the bound $\mathsf{H}(p)\leq p(1-\ln{p})$, the above probability can be bounded as
$\mathbb{P}\left[\max_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}\|\mathbf{X}_{\mathcal{A}^{c}}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{\infty}\geq 1-\frac{\bar{\alpha}}{2}\right]$ $\leq 4d\exp\left(-\frac{\bar{\alpha}^{2}n\gamma_{\min}(\mathbf{\Sigma})}{72T}+T(1+\ln(d/T))\right).$
Hence, if
$n\geq\frac{72T}{\bar{\alpha}^{2}\gamma_{\min}(\mathbf{\Sigma})}\left(T\left(1+\ln\frac{d}{T}\right)+\ln\frac{4d}{\delta}\right)=O\left(\frac{T^{2}\log(d/\delta)}{\bar{\alpha}^{2}\gamma_{\min}(\mathbf{\Sigma})}\right),$
then the above failure probability is at most $\delta$. The bound on $n$ follows from our choice of $T$. ∎
@@
The above lemma applies to standard Gaussian matrices with i.i.d. $\mathcal{N}(0,1)$ entries, in which case $\mathbf{\Sigma}=\mathbf{I}$. Hence $\bar{\alpha}=1$ and $\gamma_{\min}(\mathbf{I})=1$, and thus the mutual incoherence of a standard Gaussian matrix is greater than $1/2$ with high probability.
Finally, we analyse the last quantity, the mutual overlap between $\mathbf{y}$ and $\mathbf{X}$. The next lemma states that the mutual overlap between $\mathbf{y}$ and $\mathbf{X}$ is $\Omega(1/\sqrt{n\log{d}})$ with high probability, even if we allow the size of the set $\mathcal{A}$ to be as large as $\lfloor d/2\rfloor$.
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries and let $\mathbf{y}\in\mathbb{R}^{n}$. Let $\delta\in(0,1)$. Then
$\mathbb{P}\left[\min_{\mathcal{A}\subseteq[d]:|\mathcal{A}|<\lfloor d/2\rfloor}\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}}\geq\frac{\delta^{2/d}}{8}\sqrt{\frac{\pi}{n\ln(4nd/\delta)}}\right]\geq 1-\delta.$
@@
Let us start by bounding $\|\mathbf{X}\|_{\max}$. By a simple Chernoff bound plus a union bound, $\mathbb{P}[\|\mathbf{X}\|_{\max}\geq\sqrt{2\ln(4nd/\delta)}]\leq\delta/2$. We thus move on to the remaining part of the expression. First, for fixed $\mathcal{A}$,
$\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{\infty}=\max_{j\in\mathcal{A}^{c}}|\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}|.$
@@
Fix $j\in\mathcal{A}^{c}$. Since $\mathbf{X}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, the quantity $\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}$ is a Gaussian random variable with variance $\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{2}^{2}$. Let $\operatorname{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\text{d}t$, $z\geq 0$, be the error function. Then, for $t\in[0,1]$,
@@
$\mathbb{P}\left[\frac{|\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}|}{\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}}\leq t\sqrt{\frac{\pi}{2n}}\right]=\operatorname{erf}\left(\frac{t\sqrt{\pi}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}}{2\sqrt{n}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{2}}\right)\leq t,$
using that $\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}\leq\sqrt{n}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{2}$ and that $\operatorname{erf}(z)\leq 2z/\sqrt{\pi}$. By the independence of all $\mathbf{X}_{j}$, $j\in\mathcal{A}^{c}$, we thus have
@@
$\mathbb{P}\left[\max_{j\in\mathcal{A}^{c}}\frac{|\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}|}{\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}}\leq t\sqrt{\frac{\pi}{2n}}\right]\leq t^{d-|\mathcal{A}|}.$
Summing over all $\mathcal{A}\subseteq[d]$ such that $|\mathcal{A}|<\lfloor d/2\rfloor$, we conclude that
$\mathbb{P}\left[\min_{\mathcal{A}\subseteq[d]:|\mathcal{A}|<\lfloor d/2\rfloor}\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{\infty}}{\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{X}_{\mathcal{A}}^{\top})\mathbf{y}\|_{1}}\leq t\sqrt{\frac{\pi}{2n}}\right]\leq\sum_{|\mathcal{A}|=0}^{\lfloor d/2\rfloor-1}\binom{d}{|\mathcal{A}|}t^{d-|\mathcal{A}|}\leq 2^{d-1}t^{d/2}.$
@@
Thus, if $t=\delta^{2/d}/4$, then the above failure probability is at most $\delta/2$. By putting together both the above bound and the bound on $\|\mathbf{X}\|_{\max}$, we arrive at the desired result. ∎

### v2

@@ 4.4 Gaussian random design matrix
Define the event $\mathcal{E}(\mathbf{X}_{\mathcal{A}})\triangleq\{\|(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\|_{2}\geq 9/(n\sigma_{\min}(\mathbf{\Sigma}))\}$. Since $|\mathcal{A}|\leq n$, the event $\mathcal{E}(\mathbf{X}_{\mathcal{A}})$ happens with probability at most $2e^{-n/2}$. Therefore, with probability at least $1-2e^{-n/2}$, the quantity $\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}$ is a sub-Gaussian random variable with parameter at most $9|\mathcal{A}|/(n\sigma_{\min}(\mathbf{\Sigma}))$. A Chernoff bound yields
${\mathbb{P}}\bigg{[}{\max_{j\in\mathcal{A}^{c}}}|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\!\geq\!\frac{1-\bar{\alpha}}{2}\!\bigg{]}\!$ $\leq\!{\mathbb{P}}\bigg{[}{\max_{j\in\mathcal{A}^{c}}}|\mathbf{W}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\!\geq\!\frac{1-\bar{\alpha}}{2}\bigg{|}\mathcal{E}^{c}(\mathbf{X}_{\mathcal{A}})\!\bigg{]}\!+\!\mathbb{P}[\mathcal{E}(\mathbf{X}_{\mathcal{A}})]$
$\leq 2(d-|\mathcal{A}|)\exp\left(-\frac{(1-\bar{\alpha})^{2}n\sigma_{\min}(\mathbf{\Sigma})}{72|\mathcal{A}|}\right)+2\exp(-n/2)$
$\leq 4d\exp\left(-\frac{(1-\bar{\alpha})^{2}n\sigma_{\min}(\mathbf{\Sigma})}{72|\mathcal{A}|}\right).$
@@
The above probability is at most $\delta$ if $|\mathcal{A}|\leq(1-\bar{\alpha})^{2}\sigma_{\min}(\mathbf{\Sigma)}n/72\ln(4d/\delta)$. Thus, with probability at least $1-\delta$, $|\mathbf{X}_{j}^{\top}\mathbf{X}_{\mathcal{A}}(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{-1}\mathbf{u}|\leq\frac{1}{2}+\frac{\bar{\alpha}}{2}$ for any $\mathbf{u}\in\mathbb{R}^{|\mathcal{A}|}$ such that $\|\mathbf{u}\|_{\infty}=1$. ∎
@@
The above lemma applies to standard Gaussian matrices with $\mathbf{\Sigma}=\mathbf{I}$, in which case $\bar{\alpha}=0$ and $\sigma_{\min}(\mathbf{\Sigma})=1$, and thus the mutual incoherence $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}$ is less than $1/2$ with high probability if $|\mathcal{A}|=O(n/\log{d})$.
[…]
@@
Let $\mathbf{y}\in\mathbb{R}^{n}$ and $\mathbf{X}\in\mathbb{R}^{n\times d}$ a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble, $\mathbf{\Sigma}\succ\mathbf{0}$. Given $\mathcal{A}\subseteq[d]$, suppose that
$\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}\succeq C_{\mathbf{\Sigma}}\mathbf{I}\quad\text{for}\leavevmode\nobreak\ C_{\mathbf{\Sigma}}>0$
where $\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}=\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}^{c}}-\mathbf{\Sigma}_{\mathcal{A}^{c}\mathcal{A}}(\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}})^{-1}\mathbf{\Sigma}_{\mathcal{A}\mathcal{A}^{c}}\succ\mathbf{0}$ is the conditional covariance matrix of $\mathbf{X}_{\mathcal{A}^{c}}|\mathbf{X}_{\mathcal{A}}$. Given $\delta\in(0,1)$, if $|\mathcal{A}|\leq\min\{n/2,n-16\ln(3/\delta)\}$, then
$\mathbb{P}\left[\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\geq\frac{(\delta/3)^{2/d}}{4\|\sqrt{\mathbf{\Sigma}}\|_{1}}\frac{\|\mathbf{y}\|_{2}}{\|\mathbf{y}\|_{1}}\sqrt{\frac{\pi C_{\mathbf{\Sigma}}}{\ln(6nd/\delta)}}\right]\geq 1-\delta.$
@@
Let us start by bounding $\|\mathbf{X}\|_{\max}$. Similarly to Lemma 31, $\|\mathbf{X}\|_{\max}\leq\|\sqrt{\mathbf{\Sigma}}\|_{1}\|\mathbf{Z}\|_{\max}$, where $\mathbf{Z}$ is drawn from the $\mathbf{I}$-Gaussian ensemble. By a Chernoff bound (Fact 6) plus a union bound, $\mathbb{P}[\|\mathbf{Z}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}]\geq 1-\delta/3$, which implies that $\mathbb{P}[\|\mathbf{X}\|_{\max}\leq\sqrt{\ln(6nd/\delta)}\|\sqrt{\mathbf{\Sigma}}\|_{1}]\geq 1-\delta/3$.
We now move on to the remaining part of the expression. First notice that
$\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}=\max_{j\in\mathcal{A}^{c}}|\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}|.$
Fix $j\in\mathcal{A}^{c}$. Again by Equation 15, we can write
where $\mathbf{W}_{j}\in\mathbb{R}^{n}$ is a vector with i.i.d. $\mathcal{N}(0,[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj})$ entries. Therefore,
$|\mathbf{X}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}|=|\mathbf{W}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}|,$
@@
where we used that $\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})=\mathbf{0}$. Since $\mathbf{W}_{j}$ is independent of $\mathbf{X}_{\mathcal{A}}$, then $\mathbf{W}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}$ is a Gaussian random variable with variance
$[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{2}^{2}\geq C_{\mathbf{\Sigma}}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{2}^{2}$
since $[\mathbf{\Sigma}_{\mathcal{A}^{c}|\mathcal{A}}]_{jj}\geq C_{\mathbf{\Sigma}}$. Furthermore, since $\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+}$ is a random projection, i.e., a projection onto a random $(n-|\mathcal{A}|)$-dimensional subspace uniformly distributed in the Grassmann manifold $G_{n,n-|\mathcal{A}|}$, then Lévy’s lemma (Fact 9) yields [^4] The function $\mathbf{u}\in\mathbb{R}^{n}\mapsto\|\mathbf{P}\mathbf{u}\|_{2}$ is $1$-Lipschitz for any orthogonal projector $\mathbf{P}\in\mathbb{R}^{n\times n}$. (see also [110, Lemma 5.3.2])
$\mathbb{P}\left[\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{2}\geq\frac{1}{2}\sqrt{\frac{n-|\mathcal{A}|}{n}}\|\mathbf{y}\|_{2}\right]\geq 1-e^{-(n-|\mathcal{A}|)/16}.$
Define the event $\mathcal{E}(\mathcal{A})\triangleq\big{\{}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{2}\geq\sqrt{1-|\mathcal{A}|/n}\|\mathbf{y}\|_{2}/2\big{\}}$ and let $\operatorname{erf}(t)=\frac{2}{\sqrt{\pi}}\int_{0}^{t}e^{-z^{2}}\text{d}z$, $t\geq 0$, be the error function. Then, for $t\in[0,1]$ and conditioned on the event $\mathcal{E}(\mathcal{A})$ happening,
@@
$\mathbb{P}\left[\frac{|\mathbf{W}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}|}{\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\leq t\sqrt{\frac{\pi C_{\mathbf{\Sigma}}(n-|\mathcal{A}|)}{8n}}\frac{\|\mathbf{y}\|_{2}}{\|\mathbf{y}\|_{1}}\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \mathcal{E}(\mathcal{A})\right]\leq\operatorname{erf}\left(\frac{t\sqrt{\pi}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}{2\|\mathbf{y}\|_{1}}\right)\leq t,$
using that $\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}\leq\|\mathbf{y}\|_{1}$ and $\operatorname{erf}(t)\leq 2t/\sqrt{\pi}$. By the independence of all $\mathbf{W}_{j}$, $j\in\mathcal{A}^{c}$,
@@
$\mathbb{P}\left[\max_{j\in\mathcal{A}^{c}}\frac{|\mathbf{W}_{j}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}|}{\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\leq t\sqrt{\frac{\pi C_{\mathbf{\Sigma}}(n-|\mathcal{A}|)}{8n}}\frac{\|\mathbf{y}\|_{2}}{\|\mathbf{y}\|_{1}}\right]$ $\leq t^{d-|\mathcal{A}|}+e^{-(n-|\mathcal{A}|)/16}.$ (16)
@@
By taking $t=(\delta/3)^{1/(d-n)}$ and $|\mathcal{A}|\leq n-16\ln(3/\delta)$, then the above probability is at most $32\delta/$ 3. Finally, we put Equation 16 with $n-|\mathcal{A}|\geq n/2$ together with the bound on $\|\mathbf{X}\|_{\max}$. ∎
Remark 34.
The results from this section apply to a fixed active set $\mathcal{A}$. It is possible to bound $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$, the mutual incoherence $\alpha_{\mathcal{A}}$, and the mutual overlap $\gamma_{\mathcal{A}}$ for several sets $\mathcal{A}$ by taking a union bound on the failure probability in Lemma 31, Lemma 32, and Lemma 33. Since the approximate regularisation path $\widetilde{\mathcal{P}}$ has at most $O(n)$ kinks for when $\mathbf{X}$ is a random Gaussian matrix (Fact 22), Algorithm 3 and Algorithm 4 need only to consider $O(n)$ different active sets. Therefore, we only need to consider a union bound over $O(n)$ different sets.